# Multimodal Instruction Tuning with Conditional Mixture of LoRA

### v1

@@ Acknowledgments
This research is based upon work partially supported by the Amazon - Virginia Tech Initiative for Efficient and Robust Machine Learning.
@@ Appendix B Implementation Details
We leverage the stage-one LLaVA-v1  [^2] https://github.com/haotian-liu/LLaVA/ (before the visual instruction tuning stage) as our pre-trained large multimodal models, specifically employing LLaVA with Vicunna-7B v1.3. For all model variants, we fine-tune this stage-one LLaVa on the scale-down version of Vision-Flan for three epochs, using a total batch size of 128 and a learning rate of $54e-$ 5. The fine-tuning process for MixLoRA ($E$=16, $r$=4) takes approximately 20 hours on 4 A100 GPUs, with an effective batch size of 8 per GPU and a gradient accumulation step of 4. For LoRA, we set the hyper-paramter $\alpha$ in Equation 3.1 to be 2 $\times$ rank $r$ and for MixLoRA, we define $\alpha$ as 2 $\times$ factors $|E|$. For the other configuration, we adopt LLaVA’s default setting for LoRA fine-tuning, as provided in its codebase. For the task-specific routing, we adopt the Vicunna Chiang et al. (2023) as our pre-trained large language model $f_{\phi}(\cdot)$ for encoding task definition. Notably, Vicuna also serves as the language backbone of the LLaVA model. Following a similar approach to LoRA, for the LLaVA model with 32 Transformer layers, we insert MixLoRA into all linear layers within the Transformer layers. During training, all parameters in the MixLoRA module are updated, while the rest of LLaVA’s parameters remain frozen.

### v2

@@ Acknowledgments
This research is based upon work partially supported by the U.S. DARPA ECOLE Program #HR001122S0052, FoundSci Program #HR00112490370, and the Amazon - Virginia Tech Initiative for Efficient and Robust Machine Learning. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix B Implementation Details
We leverage the stage-one LLaVA-v1  [^3] https://github.com/haotian-liu/LLaVA/ (before the visual instruction tuning stage) as our pre-trained large multimodal models, specifically employing LLaVA with Vicunna-7B v1.3. For all model variants, we fine-tune this stage-one LLaVa on the scale-down version of Vision-Flan for three epochs, using a total batch size of 128 and a learning rate of $54e-$ 5. The fine-tuning process for MixLoRA ($E$=16, $r$=4) takes approximately 20 hours on 4 A100 GPUs, with an effective batch size of 8 per GPU and a gradient accumulation step of 4. For LoRA, we set the hyper-paramter $\alpha$ in Equation 2 to be 2 $\times$ rank $r$ and for MixLoRA, we define $\alpha$ as 2 $\times$ factors $|E|$. For the other configuration, we adopt LLaVA’s default setting for LoRA fine-tuning, as provided in its codebase. For the task-specific routing, we adopt the Vicunna Chiang et al. (2023) as our pre-trained large language model $f_{\phi}(\cdot)$ for encoding task definition. Notably, Vicuna also serves as the language backbone of the LLaVA model. Following a similar approach to LoRA, for the LLaVA model with 32 Transformer layers, we insert MixLoRA into all linear layers within the Transformer layers. During training, all parameters in the MixLoRA module are updated, while the rest of LLaVA’s parameters remain frozen.